# MIND-V: Hierarchical Video Generation for Long-Horizon Robotic Manipulation with RL-based Physical Alignment

### v1

@@ Abstract
Embodied imitation learning is constrained by the scarcity of diverse, long-horizon robotic manipulation data. Existing video generation models for this domain are limited to synthesizing short clips of simple actions and often rely on manually defined trajectories. To this end, we introduce MIND-V, a hierarchical framework designed to synthesize physically plausible and logically coherent videos of long-horizon robotic manipulation. Inspired by cognitive science, MIND-V bridges high-level reasoning with pixel-level synthesis through three core components: a Semantic Reasoning Hub (SRH) that leverages a pre-trained vision-language model for task planning; a Behavioral Semantic Bridge (BSB) that translates abstract instructions into domain-invariant representations; and a Motor Video Generator (MVG) for conditional video rendering. MIND-V employs Staged Visual Future Rollouts, a test-time optimization strategy to enhance long-horizon robustness. To align the generated videos with physical laws, we introduce a GRPO reinforcement learning post-training phase guided by a novel Physical Foresight Coherence (PFC) reward. PFC leverages the V-JEPA world model to enforce physical plausibility by aligning the predicted and actual dynamic evolutions in the feature space. MIND-V demonstrates state-of-the-art performance in long-horizon robotic manipulation video generation, establishing a scalable and controllable paradigm for embodied data synthesis.
@@ 1 Introduction
Scalable robot learning within Embodied AI [rekep, robridge, pi0] is critically bottlenecked by the scarcity of high-quality, diverse, and long-horizon robotic manipulation data [robomaster]. Video generation models [cogvideox, hunyuanvideo, wan22] offer a promising solution by potentially synthesizing an infinite stream of robotic operation videos [robomaster, robodreamer], which can fuel imitation learning and even function as world models [IRASim, wow] for embodied agents.
Figure 1: Comprehensive comparison of MIND-V against SOTA models for long-horizon robotic video generation.
@@
However, generating high-quality, long-horizon robotic manipulation videos that adhere to human commands presents significant challenges, primarily in three areas: (1) Long-Horizon Coherence Challenge: This demands maintaining causal consistency and logical flow across a sequence of interconnected sub-tasks, where a single error can compromise the entire operation [robobrain, zhai2025igniting]. (2) Semantic-to-Pixel Generation Challenge: This involves accurately translating abstract language commands into concrete spatiotemporal interactions in pixel space, which places immense demands on the model’s semantic understanding and instruction-following fidelity [controllable]. (3) Physical Plausibility Challenge: The generated output must ensure strict physical plausibility, requiring adherence to fundamental physical laws governing collision dynamics, object permanence, and interaction forces [physical_priors]. Existing methods fall short of providing a comprehensive solution to these challenges. On one hand, directly training video foundation models [hunyuanvideo, wan22, wow, cogvideox] for long-horizon tasks often suffer from logical discontinuities and detail degradation, as they struggle to bridge the vast gap from abstract commands to concrete pixel-level execution. On the other hand, while trajectory-control-based generative models [draganything, motionctrl, robomaster, IRASim] offer enhanced controllability, they do so at the cost of the autonomy and scalability essential for large-scale automated data generation.
@@
Figure 2: Overview of our hierarchical framework for long-horizon robotic manipulation video generation. Beginning in the cognitive core, the Semantic Reasoning Hub (SRH) decomposes a high-level instruction into atomic sub-tasks and plans a detailed trajectory for each. These plans are then encapsulated into our novel Behavioral Semantic Bridge (BSB), a structured, domain-invariant intermediate representation that serves as a precise blueprint for the Motor Video Generator (MVG). The MVG, a conditional diffusion model, renders photorealistic videos that strictly adhere to the kinematic constraints defined in the BSB. At inference time, Staged Visual Future Rollouts provide a “propose-verify-refine” loop for self-correction, ensuring local optimality at each stage to mitigate error accumulation.
@@
To this end, we draw inspiration from the hierarchical theory of human motor control in cognitive science [Neurobiological, Hierarchical]. The human brain executes complex tasks through a hierarchically collaborative process: high-level cognitive centers, such as the cerebral cortex, handle intent understanding and abstract planning, while low-level motor systems, like the cerebellum, translate these plans into precise muscle control. Specialized neural pathways bridge these layers, enabling the efficient translation of abstract intent into concrete physical action.
@@
Inspired by this paradigm, we introduce MIND-V, a cognition-inspired hierarchical video generation model designed to synthesize physically plausible and logically coherent long-horizon robotic manipulation videos. Emulating the brain’s cognition-to-execution pipeline, MIND-V is built upon three core components: (1) a Semantic Reasoning Hub (SRH) that performs high-level task understanding and planning based on a pre-trained Vision-Language Model (VLM); (2) a Behavioral Semantic Bridge (BSB) that acts as a task-invariant link by translating abstract plans into structured, executable representations; and (3) a Motor Video Generator (MVG) that synthesizes physically realistic manipulation videos conditioned on the BSB. By first generating the BSB through the causal reasoning of the SRH and then translating these symbolic representations into embodied actions with the MVG, MIND-V effectively bridges high-level reasoning with pixel-level synthesis through hierarchical collaboration. To the best of our knowledge, MIND-V is the first fully autonomous framework for generating long-horizon videos of embodied manipulation tasks.
@@
To further enhance physical plausibility, we design a novel RL post-training phase [marl-snake] using Group Relative Policy Optimization (GRPO) [flow_grpo], guided by a Physical Foresight Coherence (PFC) Reward. The PFC reward leverages a pre-trained world model as a “physics referee” to quantify the dynamic coherence of generated videos, thereby aligning the model with physical laws. Furthermore, to mitigate error accumulation in long-horizon tasks, we introduce the Staged Visual Future Rollouts. This test-time optimization strategy decomposes the global planning problem into a sequence of locally optimal decisions. At each sub-task transition, MIND-V performs an “propose-verify-refine” loop, where it simulates multiple future trajectories and selects the most coherent one to proceed. This staged approach effectively prevents the propagation of early-stage errors, significantly enhancing the robustness and success rate of the final video.
@@
  • We propose the first hierarchical intelligent video generation framework, MIND-V, for long-horizon robotic manipulation. MIND-V effectively bridges the gap between high-level task planning and low-level pixel synthesis through a three-tier architecture of brain (SRH), symbolic bridge (BSB), and video generator (MVG).
  • We present the Staged Visual Future Rollouts, a test-time optimization strategy that decomposes a global long-horizon generation into a series of locally optimal decisions. By performing an ’propose-verify-refine’ process at each sub-task, this method mitigates error accumulation and enhances generation robustness.
  • We propose a GRPO post-training alignment guided by a novel Physical Foresight Coherence (PFC) reward, which leverages a pre-trained world model to score the physical plausibility of generated dynamics in latent feature space, thereby steering the generator towards physically realistic outputs.
  • Experiments demonstrate that MIND-V achieves state-of-the-art performance in long-horizon robotic manipulation video generation, establishing a scalable and controllable paradigm for embodied data generation.
@@ 2 Related Work
2.1 Video Generation for Robotic Manipulation
The advancement of scalable embodied intelligence is critically dependent on large-scale, realistic data. However, collecting real-world robot data via human demonstration is a time-consuming and labor-intensive process. Video generation models [wan22, cogvideox, hunyuanvideo] have emerged as a cost-effective alternative for synthesizing photorealistic data for policy learning. Models like UniPi [UniPi] and AVDC [AVDC] frame robotic planning as a text-to-video generation problem, where imagined visual futures are subsequently translated into executable actions via inverse dynamics models. WoW [wow] and Robodreamer [robodreamer] structure video models as world models that learn latent physical dynamics from extensive interaction data to achieve compositional generalization. While these methods demonstrate strong semantic understanding, they lack fine-grained control over the precise execution of manipulation tasks [controllable]. This gap often leads to logical failures and physical inconsistencies, particularly in long-horizon scenarios. Another category of methods, such as IRASim [IRASim] and RoboMaster [robomaster], employs explicit trajectory guidance for more precise control. However, these approaches necessitate complex manual annotations, including detailed motion paths and object masks, which severely limits their scalability and autonomy. In contrast, MIND-V’s hierarchical architecture autonomously decomposes high-level commands into executable instructions for the generator, enabling the generation of long-horizon high-fidelity manipulation videos without requiring extra manual guidance.
@@ 2.2 Controllable Video Generation
Recent advancements in diffusion generation [rpd-diff, fdg-diff] have spurred demand for controllable methods that can accurately translate user intent into visual content [controllable]. To this end, research has investigated a spectrum of conditioning modalities to guide the video generation process, ranging from high-level semantic signals (text prompts [wan22]) to low-level structural inputs (masks [draganything, zhou2025fireedit, levitor, talker], trajectories [motionctrl, tora, zo3t, ma2025followyourmotion], sketches [sketchvideo, ma2024followyouremoji, VidSketch], and pose estimations [mtvcrafter, ma2024followpose, hunyuanportrait, zhang2025musetalk, flexiact, jin2024alignment]). However, these modalities present a fundamental trade-off between semantic abstraction and granular control. On the one hand, high-level semantic conditions like text can provide intuitive guidance but often falls short when dealing with complex multi-stage tasks. This abstraction can lead to semantic drift and diminished fidelity in long-horizon videos, as models struggle to map abstract language onto a consistent sequence of physical interactions [controllable]. On the other hand, low-level signals such as trajectories, masks, or sketches as conditions can offer more precise spatial and temporal control but imposes a heavy annotation burden, requiring users to specify detailed geometric constraints manually [controllable]. Our research addresses this dilemma by introducing a hierarchical framework that connects high-level reasoning with pixel-level synthesis. The Semantic Reasoning Hub (SRH) is responsible for interpreting abstract user intentions, while the Behavioral Semantic Bridge (BSB) automatically converts them into structured, multi-dimensional representations for use by the generator. This approach achieves high semantic fidelity and precise control in long-term tasks, obviating the need for auxiliary manual annotations.
@@ 3 Method
Figure 3: Architecture of the Motor Video Generator (MVG). The MVG utilizes guidance from the BSB to synthesize spatiotemporally precise videos. The process initiates with encoding the BSB’s semantic representation into the (c) Spatiotemporal Guidance Tensor, which embeds the visual features of the active agent along its planned trajectory across frames. This tensor is subsequently processed by the (b) Motion Embedding module to produce a refined motion signal ($G$). Finally, this signal is injected into the (a) Latent Diffusion Transformer, conditioning each step of the denoising process to ensure the synthesized video exhibits strict fidelity to the intended motion.
@@ 3.1 Framework
As illustrated in Figure 2, our hierarchical framework implements a top-down pipeline from high-level cognition to concrete visual representation. First, the Semantic Reasoning Hub (SRH, 3.1.1) decomposes a long-horizon task into atomic sub-tasks based on initial observations and user instructions. For each sub-task, the SRH then employs vision modules for affordance localization and trajectory planning to construct a structured, domain-invariant intermediate representation, termed the Behavioral Semantic Bridge (BSB, 3.1.2). This representation guides the Motor Video Generator (MVG,  3.1.3) in synthesizing a photorealistic video sequence. A closed-loop feedback mechanism via Staged Visual Future Rollouts returns the generated results to the SRH for evaluation and potential re-planning. A detailed description of each component follows.
3.1.1 Semantic Reasoning Hub (SRH)
@@
As the cognitive core of our framework, the SRH translates abstract semantics into actionable geometric signals. It accomplishes this by synergizing two key components: a powerful pre-trained Vision-Language Model (VLM), such as Gemini-2.5-Pro [gemini25], provides long-horizon planning and semantic reasoning, while an affordance-based visual localizer, like Affordance-R1 [affordance-r1], grounds these plans with physical common sense. This synergy transforms the SRH into a robust, physics-aware decision-making engine.
@@
Given an initial scene observation $I_{0}$ and a long-horizon task instruction $L$ (e.g., “clean the desktop”), the VLM first performs a comprehensive semantic analysis of the scene and then decomposes $L$ into an ordered sequence of atomic sub-tasks. Each sub-task is defined by a tuple $\text{SubTask}_{i}=\{\text{ActionType}_{i},\text{Object}_{i},\text{Destination}_{i}\}$, which specifies the action primitive, the object of manipulation, and its target location, respectively. This structured decomposition provides a symbolic foundation for precise downstream control.
@@
For each sub-task, the SRH leverages the affordance localizer’s vision-action alignment capabilities to precisely identify the object’s segmentation mask ($M_{\text{obj}}$) and predict its functional interaction points ($P_{obj}$, e.g., the handle of a cup). Based on this affordance data, the VLM plans a physically plausible trajectory. The trajectory is generated using a smooth curve function and discretized into a sequence of points corresponding to the video frames. This process utilizes a closed-loop refinement mechanism, where the VLM proposes a candidate trajectory that is then visualized and returned for iterative evaluation until a smooth and collision-free path is confirmed.
3.1.2 Behavioral Semantic Bridge (BSB)
@@
The BSB serves as the crucial bridge connecting high-level planning with pixel-level video synthesis. It is a structured domain-invariant intermediate representation that translates symbolic outputs from the SRH into an actionable format for the MVG. The BSB is composed of three key elements:
@@
  • Decomposed Collaborative Trajectory: For each sub-task, the trajectory is decomposed into three distinct phases: pre-interaction ($T_{\text{pre}}$, arm approaches the object), interaction ($T_{\text{interact}}$, object is manipulated), and post-interaction ($T_{\text{post}}$, arm retracts). This decomposition clearly defines the primary active agent and its objective for each stage [robomaster].
@@
By decoupling task logic from visual appearance, this design endows the BSB with domain invariance, significantly enhancing the model’s ability to generalize to novel environments and tasks.
3.1.3 Motor Video Generator (MVG)
As illustrated in Figure 3, the MVG is a conditional diffusion model built upon a Diffusion Transformer (DiT) backbone [DiT], which tasks with synthesizing manipulation videos precisely conditioned on the control signals from the BSB. To achieve this control, the MVG first encodes the BSB’s object representation into a spatiotemporal guidance tensor of size ($T\times C\times H\times W$). This tensor dynamically embeds the visual features of the active agent (arm or manipulated object) onto its planned path across the time dimension. A motion embedding module integrates this guidance into the DiT backbone during denosing process. The module employs spatiotemporal convolutions to encode the guidance tensor into a feature representation G{G}. Within each Transformer block, this representation is fused with the video’s intermediate hidden state $h$ via additive fusion:

### v2

# MIND-V: Hierarchical World Model for Long-Horizon Robotic Manipulation with RL-based Physical Alignment

@@ Abstract
Scalable embodied intelligence is constrained by the scarcity of diverse, long-horizon robotic manipulation data. Existing video world models in this domain are limited to synthesizing short clips of simple actions and often rely on manually defined trajectories. To this end, we introduce MIND-V, a cognitive hierarchical world model designed to synthesize physically plausible and logically coherent videos of long-horizon robotic manipulation. Inspired by cognitive science, MIND-V bridges high-level reasoning with pixel-level synthesis through three core components: a Semantic Reasoning Hub (SRH) that leverages a pre-trained vision-language model for task planning; a Behavioral Semantic Bridge (BSB) that translates abstract instructions into domain-invariant representations; and a Motor Video Generator (MVG) for conditional video rendering. MIND-V employs Staged Visual Future Rollouts, a test-time optimization strategy to enhance long-horizon robustness. To enforce adherence to physical laws, we introduce a GRPO reinforcement learning post-training phase guided by a novel Physical Foresight Coherence (PFC) reward. PFC leverages the V-JEPA2 world model as a “physics referee” to penalize implausible dynamics in the latent feature space. Experiments confirm MIND-V’s SOTA performance in long-horizon simulation and its significant value for policy learning, introducing a scalable and fully autonomous framework for embodied data synthesis.
@@ 1 Introduction
Scalable robot learning within Embodied AI [rekep, robridge, pi0] is critically bottlenecked by the scarcity of high-quality, diverse, and long-horizon robotic manipulation data [robomaster]. Video world models [IRASim, wow, cosmos] (VWM) that can simulate physical interactions and predict future outcomes offer a promising solution by potentially synthesizing an infinite stream of robotic operation videos [robomaster, robodreamer, dreamdojo].
Figure 1: Architectural and performance comparison of MIND-V (c) against the video foundation model (a) and trajectory-controlled generative model (b) paradigms for long-horizon robotic manipulation modeling.
@@
However, constructing a VWM capable of generating high-quality, long-horizon robotic manipulation sequences that faithfully adhere to commands presents three formidable challenges: (1) Long-Horizon Coherence Challenge: This demands maintaining causal consistency and logical flow across a sequence of interconnected subtasks, where a single error can compromise the entire operation [robobrain, zhai2025igniting]. (2) Semantic-to-Pixel Generation Challenge: This involves accurately translating abstract semantic commands or actions into concrete spatiotemporal interactions in pixel space, which places immense demands on the VWM’s semantic understanding and instruction-following fidelity [controllable]. (3) Physical Plausibility Challenge: The simulated world must ensure strict physical plausibility, requiring adherence to fundamental physical laws governing collision dynamics, object permanence, and interaction forces [physical_priors]. Existing methods fail to provide a comprehensive solution to these challenges. On one hand, directly fine-tuning video foundation models [hunyuanvideo, wan22, wow, cogvideox] for long-horizon robotic modeling often suffer from imprecise instruction following and visual degradation, as they struggle to bridge the vast gap from abstract semantic commands to concrete pixel-level execution (Fig. 1(a)). On the other hand, while trajectory-controlled generative models [draganything, motionctrl, robomaster, IRASim] offer enhanced controllability, they sacrifice the autonomy and scalability essential for large-scale automated data generation, which function more as renderers than as intelligent world simulators (Fig. 1(b)).
@@
Figure 2: Overview of MIND-V: a cognitive hierarchical video world model for long-horizon robotic manipulation. Beginning in the cognitive core, the Semantic Reasoning Hub (SRH) decomposes a high-level instruction into atomic subtasks and plans detailed trajectories for each. These plans are then encapsulated into the Behavioral Semantic Bridge (BSB), a structured, domain-invariant intermediate representation that serves as a precise blueprint for the Motor Video Generator (MVG). The MVG, a conditional diffusion model, renders photorealistic videos that strictly adhere to the kinematic constraints defined in the BSB. At inference time, Staged Visual Future Rollouts provide a “propose-verify-refine” loop for self-correction, ensuring local optimality at each stage to mitigate error accumulation.
@@
To this end, we draw inspiration from the hierarchical theory of human motor control in cognitive science [Neurobiological]. The human brain executes complex tasks through a hierarchically collaborative process: High-level cognitive centers, such as the cerebral cortex, handle intent understanding and abstract planning, while low-level motor systems, like the cerebellum, translate these plans into precise muscle control. Specialized neural pathways bridge these layers, enabling the efficient translation of abstract intent into concrete physical action [Hierarchical].
@@
Embracing this paradigm, we introduce MIND-V, a cognitive hierarchical world model designed to simulate physically plausible and logically coherent long-horizon robotic manipulations. Emulating the human brain’s “cognition-to-execution pipeline”, MIND-V is built upon three core components (Fig. 1(c)): (1) a Semantic Reasoning Hub (SRH) that performs high-level task planning based on a pre-trained Vision-Language Model (VLM); (2) a Behavioral Semantic Bridge (BSB) that acts as a domain-invariant link by translating abstract plans into structured, executable representations; and (3) a Motor Video Generator (MVG) that synthesizes physically realistic manipulation videos conditioned on the BSB. By first generating the BSB through the task reasoning of the SRH and then translating these symbolic representations into embodied action videos with the MVG, MIND-V effectively bridges high-level reasoning with pixel-level synthesis through hierarchical collaboration.
@@
To further align VWM with physical plausibility, we design a novel Group Relative Policy Optimization (GRPO) [flow_grpo] post-training [marl-snake] guided by a Physical Foresight Coherence (PFC) Reward. The PFC reward leverages a pre-trained world model as a “physics referee” to quantify the dynamic coherence of generated videos in latent feature space, thereby aligning the model with physical laws. Furthermore, to mitigate error accumulation in long-horizon tasks, we introduce the Staged Visual Future Rollouts at reference time. Based on the hierarchical architecture, this test-time optimization strategy decomposes the global planning problem into a sequence of locally optimal decisions. At each subtask transition, MIND-V performs a “propose-verify-refine” loop, where it simulates multiple future trajectories and selects the most coherent one to proceed. This staged approach effectively prevents the propagation of early-stage errors and semantic drift, significantly enhancing the fidelity and success rate of the synthesized video.
@@
  • To the best of our knowledge, MIND-V is the first hierarchical video world model for long-horizon robotic manipulation tasks. MIND-V effectively bridges the gap between high-level task planning and low-level pixel synthesis through a three-tier architecture of brain (SRH), symbolic bridge (BSB), and video generator (MVG).
  • We present the Staged Visual Future Rollouts, a test-time optimization strategy for VWM that decomposes a global long-horizon generation into a series of locally optimal decisions. By performing a “propose-verify-refine” process at each subtask, this method mitigates error accumulation and enhances generation robustness.
  • We propose a GRPO post-training guided by a novel Physical Foresight Coherence (PFC) reward, which leverages a pre-trained world model to score the physical plausibility of generated dynamics in latent feature space, thereby steering the generator towards physically realistic outputs.
  • Extensive experiments demonstrate MIND-V’s state-of-the-art performance in long-horizon robotic manipulation video synthesis and highlight its significant downstream application value, thereby establishing a scalable and controllable paradigm for robotic world modeling.
@@ 2 Related Work
2.1 Video World Models for Embodied Intelligence
The advancement of scalable embodied intelligence is critically dependent on large-scale data. However, collecting real-world robotic data is a time-consuming and labor-intensive process. Video World Models (VWM) [wan22, cogvideox, hunyuanvideo, zhang2025blockvid, team2025inferix] are capable of predicting future outcomes from current states, emerging as a efficient alternative for synthesizing photorealistic data for policy learning. Models like UniPi [UniPi] and AVDC [AVDC] frame robotic planning as a text-to-video generation problem. WoW [wow], Dreamdojo [dreamdojo] and Robodreamer [robodreamer] learn latent physical dynamics from extensive video interaction data to achieve compositional generalization. While these methods demonstrate strong semantic understanding, they lack fine-grained control over the precise execution of manipulation tasks [controllable]. This gap may lead to logical failures and physical inconsistencies, particularly in long-horizon scenarios. Another category, such as IRASim [IRASim], Cosmos [cosmos] and RoboMaster [robomaster], employs explicit trajectory guidance for more precise control and action modeling. However, these approaches necessitate complex manual annotations, including detailed motion paths and object masks, which severely limits their scalability and autonomy. In contrast, MIND-V’s hierarchical architecture autonomously decomposes high-level commands into executable instructions for the generator, enabling the generation of long-horizon high-fidelity manipulation videos without requiring extra manual guidance.
@@ 2.2 Controllable Video Generation
Recent advancements in diffusion generation [rpd-diff, fdg-diff] have spurred demand for controllable methods that can accurately translate human intent into visual content [controllable]. To this end, research has investigated a spectrum of conditioning modalities to guide the video generation process, ranging from high-level semantic signals (text [wan22] and audio [talker]) to low-level structural inputs (masks [draganything, zhou2025fireedit], trajectories [motionctrl, zo3t], sketches [ma2024followyouremoji, VidSketch], bounding boxes [liu2025controllable], and pose estimations [hunyuanportrait, zhang2025musetalk]). However, these methods present a fundamental trade-off between semantic abstraction and granular control. High-level semantic conditions like text can provide intuitive guidance but often falls short when dealing with complex multi-stage tasks. This abstraction can lead to semantic drift and diminished fidelity in long-horizon videos, as models struggle to map abstract language onto a consistent sequence of physical interactions [controllable]. On the other hand, low-level signals such as trajectories, masks, or sketches as conditions can offer more precise spatial and temporal control but imposes a heavy annotation burden, requiring users to specify detailed geometric constraints manually [controllable]. Our research addresses this dilemma by introducing a hierarchical VWM framework that connects high-level reasoning with pixel-level synthesis. The Semantic Reasoning Hub (SRH) is responsible for interpreting abstract user intentions, while the Behavioral Semantic Bridge (BSB) automatically converts them into structured, multi-dimensional representations for use by the generator. This approach achieves high semantic fidelity and precise control in long-horizon world modeling, obviating the need for auxiliary manual annotations.
@@ 3.1 Framework
As illustrated in Fig. 2, MIND-V implements a top-down pipeline from high-level cognition to concrete visual representation. First, the Semantic Reasoning Hub (SRH) decomposes a long-horizon task into atomic subtasks based on initial observations and user instructions. For each subtask, the SRH then employs vision modules for affordance localization and trajectory planning to construct a structured, domain-invariant intermediate representation, termed the Behavioral Semantic Bridge (BSB). This representation guides the Motor Video Generator (MVG) in synthesizing a photorealistic video sequence. A closed-loop feedback mechanism via Staged Visual Future Rollouts returns the generated results to the SRH for evaluation and potential re-planning. A detailed description of each component follows.
Figure 3: Architecture of the Motor Video Generator (MVG). Serving as the simulation engine, the MVG translates abstract intents (BSB) into concrete visual futures. The process initiates with encoding the BSB’s semantic representation into the (c) Spatiotemporal Guidance Tensor, which embeds the visual features of the active agent along its planned trajectory. This guidance is then processed by the (b) Motion Embedding module to produce refined motion signals ($G$). These signal are finally injected into the (a) Latent Diffusion Transformer, conditioning the denoising process to ensure the synthesized video strictly adheres to the intended physical dynamics.
@@
Semantic Reasoning Hub (SRH) As the cognitive core of our VWM, the SRH translates abstract semantics into actionable geometric signals. It accomplishes this by synergizing two key components: a powerful pre-trained Vision-Language Model (VLM), such as Gemini-2.5-Pro [gemini25], provides long-horizon planning and semantic reasoning, while an affordance-based visual localizer, like Affordance-R1 [affordance-r1], grounds these plans with physical common sense. This synergy transforms the SRH into a robust, physics-aware decision-making engine.
@@
Given an initial scene observation $I_{0}$ and a long-horizon task instruction $L$ (e.g., “clean the desktop”), the VLM first performs a comprehensive semantic analysis of the scene and then decomposes $L$ into an ordered sequence of atomic subtasks. Each subtask is defined by a tuple $\text{SubTask}_{i}=\{\text{ActionType}_{i},\text{Object}_{i},\\ \text{Destination}_{i}\}$, which specifies the action primitive, the object of manipulation, and its target location, respectively. This structured decomposition provides a symbolic foundation for precise downstream control generation.
@@
For each subtask, the SRH leverages the affordance localizer’s vision-action alignment capabilities to precisely identify the object’s segmentation mask ($M_{\text{obj}}$) and predict its functional interaction points ($P_{obj}$, e.g., the handle of a cup). Based on this affordance data, the VLM plans a physically plausible trajectory. The trajectory is generated using a smooth curve function and discretized into a sequence of points corresponding to the video frames.
@@
Behavioral Semantic Bridge (BSB)  The BSB serves as the crucial bridge connecting high-level planning with pixel-level video synthesis. It is a structured domain-invariant intermediate representation that translates symbolic outputs from the SRH into an actionable format for the MVG. The BSB is composed of three key elements:
@@
  • Decomposed Collaborative Trajectory: For each subtask, the trajectory is decomposed into three distinct phases: pre-interaction ($T_{\text{pre}}$, arm approaches the object), interaction ($T_{\text{interact}}$, object is manipulated), and post-interaction ($T_{\text{post}}$, arm retracts). This decomposition clearly defines the primary active agent and its objective for each stage [robomaster].
@@
By decoupling task logic from visual appearance, this design endows the BSB with domain invariance, significantly enhancing the our VWM’s ability to generalize to novel environments and tasks.
Motor Video Generator (MVG)  As depicted in Fig. 3, the MVG functions as a learnable physics engine. Built upon a Diffusion Transformer (DiT) backbone [DiT], it is trained to synthesize high-fidelity video sequences that strictly adhere to the kinematic constraints defined by the BSB. To enforce this control, the MVG first encodes the BSB’s object representation into a spatiotemporal guidance tensor of size ($T\times C\times H\times W$). This tensor dynamically embeds the visual features of the active agent (arm or manipulated object) onto its planned path across the time dimension. A motion embedding module integrates this guidance into the DiT backbone during denoising process. The module employs spatiotemporal convolutions to encode the guidance tensor into a feature representation G{G}. Within each DiT block, this representation is fused with the video’s intermediate hidden state $h$ via additive fusion: